A hydraulic lift has pistons that are 1 cm and 15 cm in diameter. If a...

  1. A hydraulic lift has pistons that are 1 cm and 15 cm in diameter. If a force of 180 N is applied to the master cylinder, what is the largest mass that the lift can support?
Answer #1

Answer: According to Bernoulli's principle, the pressure at a point in a static fluid is transmitted equally throughout the fluid undiminished.

Now, in a hydraulic lift, we use this to our advantage by placing two cylinders of different cross-sectional areas on a static fluid. Then we apply force f on the smaller cylinder of area a, which gives pressure P=f/a.

Since the pressure is constant in the static fluid, the pressure reverted back on the large cylinder will incur a higher force, since

P=F/A=f/a

So, F=f/a * A = 180/(*1*1) * (*15*15) = 40500 N

Now, the 40500 N of force can only sustain/lift a mass whose weight is less than equal to 40500 N

So, the maximum mass possible is 40500/g = 40500/9.8 = 4132.65 Kg